Pictured on the right are thee point charges Q_1 =
Answer #1

Potential due to a point charge q at distance r is V = Kq/d

And if there are multiple charges, the net potential at a point is scalar sum of potentials due to individual charges.

let a is the side of square , a = 40.1 cm = 0.401 m

VA = KQ1/a +KQ2/a*1.414 + KQ3/a

VA =9*109*[18.4*10-6/0.401 - 36.6*10-6/ (1.414*0.401) + 57.3*10-6/0.401]

V = 1118.152*103 v